1.)calculate the volume, in mL of 1.72M HCl solution that will react with 2.67 mol of CaCO3? (Ans. 3.10x103mL



___HCl+___CaCO3____CaCl2+___H2O+__CO2



2.)If 0.650% solution of NaCO3 will be used to precipitate Ca+2 from solution, determine the amount of solution, in grams, needed to precipitate 5.00x102mL of 0.01120M Ca+2?(Atomic mass :Na=22.99g/mol:C=12.01g/mol:O=16.00g/mol:Ca=40.08g/mol:Ans.91.3g)



____Na2CO3(aq)+___Ca+2(aq)____CaCO3(s)+___Na+(aq)

1
Expert's answer
2022-05-30T17:01:10-0400

1).

CaCO3 + 2 HCl = CaCl2 + CO2 + H2O

According to the above reaction,

1 mol CaCO3 —> 2 mol HCl

2,76mol CaCO3 —> X

X = 2,76 * 2 / 1 = 5,52 mol HCl


V = n / CM = 5,52mol / 1,72mol/l = 3,21 litr or 3,21 * 103 ml

V — volume of solution (litr);

n — amount of substance (mol);

CM — concentration of molarity.

ANSWER: 3,21 * 103 ml.


2).

a)

V — volume = 5*102ml = 0,5 litr;

CM — concentration of molarity =

= 0,0112 mol/l;

n — amount of substance (mol) = ?


n (Ca2+) = V * CM = 0,5 * 0,0112 =0,0056 mol

b)

Mr (Na2CO3) = 23*2+12+16*3 = 106 g/mol

Na2CO3 + Ca2+ => CaCO3 + 2Na+

According to the above reaction,

106 g Na2CO3 — > 1mol Ca2+

X — > 0,0056mol Ca2+

X = 106 * 0,0056 / 1 = 0,5936 g Na2CO3

c)

0,5936 g — 0,65%

X — 100%

X = 0,5936 * 100 / 0,65 = 91,3 g

ANSWER: 91,3 g